OpenStudy (anonymous):

suppose that f(x) = g(h(x) ). Find possible formulas for g(x) and h(x) (There may be more than one possible answer. Assume g(x) does not equal x and h(x) does not equal x. f(x) = 1/x2 + 8x + 16

zepdrix (zepdrix):

\[g(x)=\frac{ 1 }{ x }\]\[h(x)=x^2+8x+16\] Here is one possibility. Understand how I did that? We'll end up sticking the function h(x) into the x slot of g(x).

zepdrix (zepdrix):

\[f(x)=\frac{ 1 }{ x^2+8x+16 }\] Hopefully I was reading your f(x) correctly. It should look like this right? :o

OpenStudy (anonymous):

can you explain how to decide what goes with h(x) and which one goes with g(x)? is there some sort of formula or equation I can use to help me understand how formulas are derived.

zepdrix (zepdrix):

hmm

zepdrix (zepdrix):

 I'm not sure that this answers your question :( But it's just another example. Another way to think about the ARGUMENT of a function. x being the default argument in g(x).